# On the Sylvester program and Cayley algorithm for vector partition reduction

### v3

@@ 1 Introduction
Sylvester proved [17] a statement about splitting of the scalar partition into periodic and non-periodic parts and showed that the partition function may be presented as a sum of ”waves”
@@
The first step in this direction was made in the 19th century by J.J. Sylvester [16, 17] who suggested an iterative procedure of reduction of a vector partition into a sum of scalar partitions; in [16] he gave an example of such a reduction for a double partition. Based on this approach A. Cayley developed an algorithm of double partition reduction and established a set of conditions of this method applicability [4]. Cayley showed that each column ${\bf c}_{i}$ of the two-row matrix ${\bf D}$ gives rise to a scalar partition and elements of this column must be relatively prime (we call it P-column). Another limitation of the Cayley method is that the columns ${\bf c}_{i}$ should represent noncollinear vectors. An application of the Cayley algorithm to the computation of the Gaussian polynomial coefficients is considered in [13]. A short historical excurse into the (seemingly forgotten) ideas of Sylvester and Cayley is given in Section 2 while the essence of the original Cayley algorithm is presented in Section 3.1.
The author of this manuscript considered the case when the Cayley formula for double partition fails and discussed its modification that successfully resolves the problem [14]. Specifically, the elimination of a column with the greatest common divisor (GCD) of its elements $\mbox{gcd}({\bf c}_{i})>1$ (we call it NP-column) leads not to a single scalar partition but to a weighted sum of such partitions with modified argument. This result is described in Section 3.2 and discussed in Section 3.3.
In Section 4 we extend the reduction algorithm to $l$-tuple (for arbitrary $l\geq 2$) vector partition defined by a $l\times m$ nonnegative generator matrix ${\bf D}$ made of $m$ columns ${\bf c}_{i}$. The first step of the algorithm discussed in Section 4.1 is the partial fraction expansion of the generating function (6) into $m$ terms corresponding to elimination of each of $m$ unknowns. The contribution of a column is computed in Section 4.2 and it is either a single partition for P-columns (Section 4.2.1) or a superposition of such partitions (Section 4.2.2) for NP-column. When the last element of a column is zero both the original Cayley method and its generalization cannot be applied directly so that this case requires a special treatment. In Section 4.3 we prove that the contribution of such a column to the vector partition is zero.
The case of collinear columns of the generator matrix ${\bf D}$ is addressed in Section 5 where we show that the vector partition reduces to a convolution of scalar partitions. Additional details of the iterative reduction procedure are discussed in Section 6 while the concluding remarks are presented in Section 7.
@@ 2 Sylvester-Cayley method
Then he added in the footnote: ”If there be $r$ simultaneous simple equations between $n$ variables (in which the coefficients are all positive or negative integers) forming a definite system (that is, one in which no variable can become indefinitely great in the positive direction without one or more of the others becoming negative), and if the $r$ coefficients belonging to each of the same variable are exempt from a factor common to them all, and if not more than $r-1$ of the variables can be eliminated simultaneously between the $r$ equations, then the determination of the number of positive integer solutions of the given system may be made to depend on like determinations for each of $n$ derived independent systems, in each of which the number of variables and equations is one less than in the original system” [16].
@@
In other words, Sylvester claimed that $r$-tuple vector partition for $n$ variables can be reduced to a sum of $\binom{n}{r-1}$ scalar partitions and the reduction algorithm is an iterative process based on the variable elimination. Sylvester considered a specific double partition problem as an illustration of his method and determined regions (chambers) on a plane each having a unique expression for vector partition valid in this region only. He showed that the expressions in the adjacent chambers coincide at their common boundary (see also [15]).
@@
Two years later this approach was successfully applied by Cayley [4] to double partitions subject to some restrictions on the positive elements of the matrix ${\bf D}$ (the columns ${\bf c}_{i}$ are noncollinear, column elements should be relatively prime and for all $1\leq i\leq m$ the inequality $c_{i2}<s_{2}+2$ holds). Cayley noticed : ”The subject (as I am aware) has hardly been considered except by Professor Sylvester, and it is greatly been regretted that only an outline of his valuable researches has been published: the present paper contains the demonstration of a theorem, due to him, by which (subject to certain restrictions) the question of Double Partitions is made to depend upon the ordinary theory of Single Partitions” [4]. Cayley demonstrated that row elimination leads to a sum of $m$ scalar partition terms each corresponding to elimination of single column of the matrix ${\bf D}$.
@@
It should be noted that both Sylvester and Cayley usually considered coefficients of the system of Diophantine equations ${\bf s}={\bf D}\cdot{\bf x}$ being positive and subject to the specific conditions, while once Cayley mentioned that the elements of the matrix ${\bf D}$ as well as $s_{i}$ “being all positive integer numbers, not excluding zero” [4] but never explicitly discussed such a case.
@@
Both manuscripts [16] and [4] were mentioned by L.E. Dickson in the then comprehensive volume [6] on the theory of numbers devoted to Diophantine analysis, but the literature search shows that both the reduction idea of Sylvester and the result of Cayley were forgotten. There are no publications discussing Sylvester research program of vector partition reduction to scalar ones. In my opinion the ideas and results of Sylvester and Cayley deserve not to be neglected, and this manuscript is an attempt to revive them and pave way to a general algorithm for vector partition expression through scalar partitions only.
@@ 3 Cayley method for double partitions
Computation of vector partition $W({\bf s},{\bf D})$ in (7) is performed by iterative elimination of $(l-1)$ variables $t_{k},(2\leq k\leq l)$ [2]. Each elimination step includes partial fraction expansion (PFE) w.r.t. the eliminated variable with subsequent coefficient evaluation. This step is equivalent to elimination of $k$-th row of augmented $l\times(m+1)$ matrix ${\bf E}=\{{\bf s}={\bf c}_{0},{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}$ made of the matrix ${\bf D}$ and vector ${\bf s}$; the same time the numerical columns of ${\bf D}$ are eliminated one by one. The number of new $(l-1)\times m$ matrices is equal to $m$. This algorithm was employed in [4] for a two-row positive matrix (see also [16]). The original Cayley algorithm and its generalization are exhaustively discussed in [14] and below we present their main points.
@@ 3.1 Original Cayley algorithm
[…]
@@ 3.2 Cayley method generalization
Introduce $(m-1)$-dimensional vectors
@@
with $10\leq k_{i}\leq\beta_{1}-$ 1. Expanding the r.h.s. of (18) we obtain
@@ 3.3 Cayley method for P- and NP-columns
Note that for $\beta_{1}=0$ the expression for $S_{1}(x,y_{0})$ in (18) vanishes implying vanishing contribution $C^{1}_{r,\rho}=0$. We also observe that the result (15) with $i=1$ and $\beta_{1}=0$ formally produces all $d_{1j}<0$ and $L_{1}({\bf s})=-s_{2}b_{1}$ leading to negative $L^{\prime}_{1}({\bf s})=-s_{2}b_{1}+\sum_{j=2}^{m}d_{1j}$ for any positive $s_{2}$. This means that the scalar partition $W(L_{1}({\bf s}),{\bf d}_{1})=W(L^{\prime}_{1}({\bf s}),|{\bf d}_{1}|)$ equals zero. Thus the rule of thumb for double partitions with nonnegative generators $b_{i},\beta_{i}$ and $r,\rho\geq 0$ is – the contribution of the column with $\beta_{i}=0$ is zero.
The same time we have to recall that in the course of the derivation of (15) we explicitly assumed that $\beta_{1}\neq 0$. Nevertheless it is possible to show that in case $\beta_{1}=0$ the contribution $C^{1}_{r,\rho}$ indeed vanishes. It appears to be a particular case of the general result valid for a $l$-tuple partition ($l\geq 2$) with the zero last element $c_{1l}=0$ in the first column ${\bf c}_{1}$ of the $l\times m$ matrix ${\bf D}$. The reader is referred to Section 4.3 where the details of the derivation are presented.
@@ 4.1 Partial fraction expansion
[…]
@@
[…]
@@
where the rational exponent $\nu_{k}$ of $t_{k}$ is split into an integer $j_{k}$ and a fractional part. The fractional part of $\nu_{k}$ gives an exponent $j_{l}$ of $t_{l0}$, while $t_{k}^{j_{k}}$ contributes to $A_{1,j_{l}}(\hat{\bf t})$ in (31) as a factor in $\hat{\bf t}^{\hat{\bf j}}$. We observe that
@@
[…]
@@ 4.2.1 Compact form
$C_{\bf s}^{1}=C[\hat{\bf t}^{\hat{\bf s}-s_{l}\hat{\bf c}_{1}/c_{1l}}]\left(A_{1}(\hat{\bf t},\hat{\bf t}^{-\hat{\bf c}_{1}/c_{1l}})\right)=C[\hat{\bf t}^{c_{1l}\hat{\bf s}-s_{l}\hat{\bf c}_{1}}]\left(A_{1}(\hat{\bf t}^{c_{1l}},\hat{\bf t}^{-\hat{\bf c}_{1}})\right)=C[\hat{\bf t}^{c_{1l}\hat{\bf s}-s_{l}\hat{\bf c}_{1}}]\left(1/\Pi_{1}(\hat{\bf t})\right),$ (44)
@@
where $\tilde{\bf c}^{1}_{i}$ for the columns not affected by the procedure is equal to its original value $\tilde{\bf c}^{1}_{i}={\bf c}^{1}_{i}$. The factor $H(\tilde{s}^{1}_{l-1})$ reflects the fact that the $(l-1)$-th element of the column $\tilde{\bf s}^{1}$ should be nonnegative $\tilde{s}^{1}_{l-1}\geq 0$.
Note that (43) does not work when GCD of the column ${\bf c}_{1}$ elements is larger than unity: $\gcd({\bf c}_{1})\equiv\gcd(c_{11},c_{12},\ldots,c_{1l})=g_{1}>1$. Then we have for such NP-column ${\bf c}_{1}=g_{1}{\bf c}^{\star}$ leading to $\hat{\bf c}_{1}/c_{1l}=\hat{\bf c}^{\star}/c_{l}^{\star}$ and $t_{l0}^{j_{l}}=\hat{\bf t}^{-j_{l}\hat{\bf c}_{1}/c_{1l}}=\hat{\bf t}^{-j_{l}\hat{\bf c}^{\star}/c_{l}^{\star}}$ enters the sum in (43) where $11\leq j_{l}\leq c_{1l}-$ 1, but $c_{1l}-1>c_{l}^{\star}$. The summation does not evaluate into $S_{1}(\hat{\bf t},t_{l0})$ and thus both (43) and (45) fail. The case of NP-column requires an approach that leads to an expanded form of $C_{\bf s}^{1}$ contribution.
@@ 4.2.2 General form
4.3 Zero in last row of matrix ${\bf D}$
Consider evaluation of a $l$-tuple partition with $c_{1l}=0$. Rewrite the system ${\bf s}={\bf D\cdot x}$ as
@@ 4.3 Zero in last row of matrix ${\bf D}$
[…]
@@
[…]
@@
If the matrix $\bm{D}^{\prime}_{1}$ has another column ${\bf c}_{j}$ with last zero element we have two options – either the columns ${\bf c}_{1}$ and ${\bf c}_{j}$ are collinear or they are not. In the case of noncollinear columns the contributions $C_{\bf s}^{i}$ are independent of each other and thus we have $C_{\bf s}^{1}=C_{\bf s}^{j}=0$ while for the collinear columns one should use the method described in Section 5.
5 Vector partition with collinear columns
The reduction algorithm described in Section 4 cannot be applied in case when in the generator matrix ${\bf D}$ has collinear columns as it leads to appearance of the zero columns in the reduced partition.
Call an integer generator matrix competent if its columns (viewed as vectors) are noncollinear and incompetent otherwise. Assume that the vectors corresponding to the first $n$ columns of the incompetent generator matrix ${\bf D}$ are parallel to each other and rewrite the linear system ${\bf D}\cdot{\bf x}={\bf s}$ as follows

### v4

@@ 1 Introduction
The first attempt to consider the analytical properties of the function $W(s,{\bf d})$ was made by Cayley in his pioneering work [4] where he considered the partial fraction decomposition of the generating function $G(t,{\bf d})$. Cayley found that the scalar partition can be written as a sum of a polynomial and expressions of the other type containing the factors with integer periods. Based on these findings Sylvester proved [18] a theorem about splitting of the scalar partition into periodic and non-periodic parts and showed that the partition function may be presented as a sum of ”waves”
@@
The first step in this direction was made in the 19th century by J.J. Sylvester [19, 18] who suggested an iterative procedure of reduction of a vector partition into a sum of scalar partitions; in [19] he gave an example of such a reduction for a double partition. The method is based on the elimination one-by-one of the variables $x_{i}$ from the linear system ${\bf s}={\bf D}\cdot{\bf x}$. Based on this approach A. Cayley developed an algorithm of double partition reduction and established a set of conditions of the method applicability [5]. Cayley showed that each column ${\bf c}_{i}$ of the two-row matrix $\bf D$} gives rise to a scalar partition and elements of this column must be relatively prime (we call it P-column). Another limitation of the Cayley method is that the columns ${\bf c}_{i}$ should represent noncollinear vectors. An application of the Cayley algorithm to the computation of the Gaussian polynomial coefficients is considered in [14]. A short historical excurse into the (seemingly forgotten) ideas of Sylvester and Cayley is given in Section 2 while the essence of the original Cayley algorithm is presented in Section 3.1.
The author of this manuscript considered the case when the Cayley formula for double partition fails and discussed its modification that successfully resolves the problem [15]. Specifically, the elimination of a column with the greatest common divisor (GCD) of its elements $\mbox{gcd}({\bf c}_{i})>1$ (NP-column) leads not to a single scalar partition but to a weighted sum of such partitions with modified argument. This result is described in Section 3.2 and discussed in Section 3.3.
In Section 4 we extend the reduction algorithm to $l$-tuple (for arbitrary $l\geq 2$) vector partition defined by a $l\times m$ nonnegative generator matrix $\bf D$} made of $m$ columns ${\bf c}_{i}$. The first step of the algorithm discussed in Section 4.1 is the partial fraction expansion of the generating function (6) into $m$ terms corresponding to elimination of each of $m$ unknowns. The contribution of a column is computed in Section 4.2 and it is either a single partition for P-columns (Section 4.2.1) or a superposition of such partitions (Section 4.2.2) for NP-column. When the last element of a column is zero (LZ-column) both the original Cayley method and its generalization cannot be applied directly so that this case requires a special treatment. In Section 4.3 we prove that the contribution of such a column to the vector partition is zero.
The case of collinear columns (C-columns) of the generator matrix $\bf D$} is first addressed in Section 5.1 where we show that the vector partition reduces to a convolution of scalar partitions. An alternative approach of dealing with the C- and NP-columns was suggested by Sylvester [20], namely, to enlarge the system ${\bf s}={\bf D}\cdot{\bf x}$ to eliminate these columns and then use the variable elimination procedure of partition reduction. The application of this method is discussed in Section 5.2. Additional details of the iterative reduction procedure are discussed in Section 6 while the concluding remarks are presented in Section 7.
@@ 2 Sylvester-Cayley method
Then he added in the footnote: ”If there be $r$ simultaneous simple equations between $n$ variables (in which the coefficients are all positive or negative integers) forming a definite system (that is, one in which no variable can become indefinitely great in the positive direction without one or more of the others becoming negative), and if the $r$ coefficients belonging to each of the same variable are exempt from a factor common to them all, and if not more than r1r-1 of the variables can be eliminated simultaneously between the $r$ equations, then the determination of the number of positive integer solutions of the given system may be made to depend on like determinations for each of $n$ derived independent systems, in each of which the number of variables and equations is one less than in the original system” [19].
@@
In other words, Sylvester claimed that $r$-tuple vector partition for $n$ variables can be reduced to a sum of $\binom{n}{r-1}$ scalar partitions and the reduction algorithm is an iterative process based on the variable elimination. Sylvester also described the general conditions of the method applicability – each column should be the coprime set, and the elimination of any variable should not lead to the elimination of any other variable, i.e., the columns should be noncollinear. Sylvester considered a specific double partition problem as an illustration of his method and determined regions (chambers) on a plane each having a unique expression for vector partition valid in this region only. He showed that the expressions in the adjacent chambers coincide at their common boundary (see also [17]).
@@
Two years later this approach was successfully applied by Cayley [5] to double partitions subject to the Sylvester conditions on the matrix $\bf D$} elements. with the added requirement that the inequalities $c_{i2}<s_{2}+2,\ 1\leq i\leq m,$ should hold. In the opening paragraph Cayley noticed: ”The subject (as I am aware) has hardly been considered except by Professor Sylvester, and it is greatly been regretted that only an outline of his valuable researches has been published: the present paper contains the demonstration of a theorem, due to him, by which (subject to certain restrictions) the question of Double Partitions is made to depend upon the ordinary theory of Single Partitions” [5]. The manuscript demonstrated that the procedure leads to a sum of $m$ scalar partition terms each corresponding to elimination of single variable $x_{i}$ from the vector $\bf x$}.
@@
It should be noted that both Sylvester and Cayley usually considered coefficients of the system of Diophantine equations ${\bf s}={\bf D}\cdot{\bf x}$ being positive and only once Cayley mentioned that the elements of the matrix $\bf D$} as well as $s_{i}$ “being all positive integer numbers, not excluding zero” [5] but never explicitly discussed such a case.
@@
Both manuscripts [19] and [5] were mentioned by L.E. Dickson in the then comprehensive volume [7] on the theory of numbers devoted to Diophantine analysis, but the literature search shows that both the reduction idea of Sylvester and the result of Cayley were forgotten. There are no publications discussing Sylvester research program on vector partition reduction to scalar ones. In my opinion the ideas and results of Sylvester and Cayley deserve not to be neglected, and this manuscript is an attempt to revive them and pave way to a general algorithm for vector partition expression through scalar partitions only.
@@ 3 Cayley method for double partitions
Computation of vector partition $W({\bf s},{\bf D})$ in (7) is performed by iterative elimination of (l1)(l-1) variables $t_{k},(2\leq k\leq l)$ [2]. Each elimination step includes partial fraction expansion (PFE) w.r.t. the eliminated variable with subsequent coefficient evaluation. This step is equivalent to elimination of $k$-th row of augmented $l\times(m+1)$ matrix ${\bf E}=\{{\bf s}={\bf c}_{0},{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}$ made of the matrix $\bf D$} and vector $\bf s$}; the same time the numerical columns of $\bf D$} are eliminated one by one. The number of new $(l-1)\times m$ matrices is equal to $m$. This algorithm was employed in [5] for a two-row positive matrix (see also [19]). The original Cayley algorithm and its generalization are exhaustively discussed in [15] and below we present their main points.
@@ 3.1 Original Cayley algorithm
Repeating computation for each $T_{i}(x,y),\ 1\leq i\leq m,$ we obtain for double partition
@@ 3.2 Cayley method generalization
Introduce (m1)(m-1)-dimensional vectors
@@
with $0\leq k_{i}\leq\beta_{1}-1$. Expanding the r.h.s. of (18) we obtain
@@ 3.3 Cayley method for P- and NP-columns
Note that for $\beta_{1}=0$ the expression for $S_{1}(x,y_{0})$ in (18) vanishes implying vanishing contribution $C^{1}_{r,\rho}=0$. We also observe that the result (15) with $i=1$ and $\beta_{1}=0$ formally produces all $d_{1j}<0$ and $L_{1}({\bf s})=-s_{2}b_{1}$ leading to negative $L^{\prime}_{1}({\bf s})=-s_{2}b_{1}+\sum_{j=2}^{m}d_{1j}$ for any positive $s_{2}$. This means that the scalar partition $W(L_{1}({\bf s}),{\bf d}_{1})=W(L^{\prime}_{1}({\bf s}),|{\bf d}_{1}|)$ equals zero. Thus the rule of thumb for double partitions with nonnegative generators $b_{i},\beta_{i}$ and $r,\rho\geq 0$ is – the contribution of the LZ-column with the last zero $\beta_{i}=0$ is zero.
The same time we have to recall that in the course of the derivation of (15) we explicitly assumed that $\beta_{1}\neq 0$. Nevertheless it is possible to show that in case $\beta_{1}=0$ the contribution $C^{1}_{r,\rho}$ indeed vanishes. It appears to be a particular case of the general result valid for a $l$-tuple partition ($l\geq 2$) the first LZ-column ${\bf c}_{1}$ of the $l\times m$ matrix $\bf D$}. The reader is referred to Section 4.3 where the details of the derivation are presented.
@@ 4.1 Partial fraction expansion
where functions $A_{i}$ are rational in $\hat{\bf t}$ and rational and integral (of degree $c_{il}-1$) in $t_{l}$. Consider the term $A_{1}(\hat{\bf t},t_{l})$ that corresponds to elimination of the last row and the first column ${\bf c}_{1}$ of $\bf E$}. Introduce matrix ${\bf E}_{1}$ produced from $\bf E$} by dropping the column ${\bf c}_{1}$ and present it as an array of $l$ rows ${\bf r}_{1k},\ 1\leq k\leq l,$ each having $m$ elements ${\bf E}_{1}=\{{\bf r}_{11},{\bf r}_{12},\ldots,{\bf r}_{1l}\}$. The term $A_{1}(\hat{\bf t},t_{l})$ for $t_{l0}=\hat{\bf t}^{-\hat{\bf c}_{1}/c_{1l}}$ turns into
@@
where $A_{1,j_{l}}$ are rational functions in $\hat{\bf t}$. Define $l$ vectors $\bm{r}^{\prime}_{1k},\ 1\leq k\leq l,$ by retaining last (m1)(m-1) (only numerical) elements of vectors ${\bf r}_{1k}$, and a set of $c_{1l}^{m-1}$ vectors $\bm{K}^{\prime}_{p}=\{k_{2},k_{3},\ldots,k_{m}\}$ with $0\leq k_{i}\leq c_{1l}-1$. Expanding the product in (31) we obtain a sum
@@
[…]
@@
$\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l}=c_{1l}t+j_{l},\ t=\lfloor(\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{1l})/c_{1l}\rfloor,\ 0\leq t=T_{1}(j_{l})\leq\lfloor(R_{1}-j_{l})/c_{1l}\rfloor,\ R_{1}=(c_{1l}-1)|\bm{r}^{\prime}_{1l}|.$ (34)
@@ 4.2.1 Compact form
[…]
@@
where $\tilde{\bf c}^{1}_{i}$ for the columns not affected by the procedure is equal to its original value $\tilde{\bf c}^{1}_{i}={\bf c}^{1}_{i}$. The factor $H(\tilde{s}^{1}_{l-1})$ reflects the fact that the (l1)(l-1)-th element of the column $\tilde{\bf s}^{1}$ should be nonnegative $\tilde{s}^{1}_{l-1}\geq 0$.
Note that (43) does not work when GCD of the column ${\bf c}_{1}$ elements is larger than unity: $\gcd({\bf c}_{1})\equiv\gcd(c_{11},c_{12},\ldots,c_{1l})=g_{1}>1$. Then we have for such NP-column ${\bf c}_{1}=g_{1}{\bf c}^{\star}$ leading to $\hat{\bf c}_{1}/c_{1l}=\hat{\bf c}^{\star}/c_{l}^{\star}$ and $t_{l0}^{j_{l}}=\hat{\bf t}^{-j_{l}\hat{\bf c}_{1}/c_{1l}}=\hat{\bf t}^{-j_{l}\hat{\bf c}^{\star}/c_{l}^{\star}}$ enters the sum in (43) where $1\leq j_{l}\leq c_{1l}-1$, but $c_{1l}-1>c_{l}^{\star}$. The summation does not evaluate into $S_{1}(\hat{\bf t},t_{l0})$ and thus both (43) and (45) fail. The case of NP-column requires an approach that leads to an expanded form of $C_{\bf s}^{1}$ contribution.
@@ 4.2.2 General form
4.3 Vector partitions with LZ-columns
Consider evaluation of a vector partition having the first LZ-column with $c_{1l}=0$. Rewrite the system ${\bf s}={\bf D}\cdot{\bf x}$ as
@@ 4.3 Vector partitions with LZ-columns
Introduce $\bm{x}^{\prime}_{1}=\{x_{2},x_{3},\ldots,x_{m}\}$ and use the vectors $\bm{r}^{\prime}_{1k},\ 1\leq k\leq l,$ defined in Section 4.1 to obtain a system of $l$ equations
@@
where we assume that the last row of $\bm{D}^{\prime}_{1}$ is free of zeros. The relations (45) determine the contribution $W_{i}^{l}({\bf s})$ obtained by elimination of the P-column ${\bf c}_{i},\ i>1$ as $W({\bf s}^{i},{\bf D}_{i})$ with ${\bf D}_{i}=\{{\bf c}^{i}_{1},{\bf c}^{i}_{2},\ldots,{\bf c}^{i}_{m}\}$ and ${\bf c}^{i}_{j}=c_{il}\hat{\bf c}_{j}-c_{jl}\hat{\bf c}_{i},\ 1\leq j\neq i\leq m,$ where $\hat{\bf c}_{i}$ denotes the column ${\bf c}_{i}$ without its last element $c_{il}$. The explicit expression for ${\bf c}^{i}_{1}$ reads ${\bf c}^{i}_{1}=c_{il}\hat{\bf c}_{1}-c_{1l}\hat{\bf c}_{i}=c_{il}\hat{\bf c}_{1}$. The column ${\bf s}^{i}$ is evaluated to ${\bf s}^{i}=c_{il}\hat{\bf s}-s_{l}\hat{\bf c}_{i}$.
@@
If the matrix $\bm{D}^{\prime}_{1}$ has another LZ-column ${\bf c}_{j}$ we have two options – either the columns ${\bf c}_{1}$ and ${\bf c}_{j}$ are collinear or they are not. In the case of noncollinear LZ-columns the contributions $C_{\bf s}^{i}$ are independent of each other and thus we have $C_{\bf s}^{1}=C_{\bf s}^{j}=0$. Application of the Cayley reduction algorithm (or its extended version) to the partition $W({\bf s},{\bf D})$ with two collinear LZ-columns ${\bf c}_{i}$ and ${\bf c}_{j}$ produces two partitions $W({\bf s}^{i},{\bf D}_{i})$ and $W({\bf s}^{j},{\bf D}_{j})$ where both matrices ${\bf D}_{i},{\bf D}_{j}$ have one column made of only zeros. Generally such a partition would correspond to the infinite number of solutions of the system ${\bf s}^{i}={\bf D}_{i}\cdot{\bf x}$, but it is not true in this specific case. As the vector ${\bf s}^{i}$ has at least one negative element the corresponding system has no solutions and thus $C_{\bf s}^{i}=C_{\bf s}^{j}=0$.
5 Vector partitions with C-columns
The reduction algorithms described in Sections 3,4 cannot be applied in case when in the generator matrix $\bf D$} has the collinear columns (C-columns) as it leads to appearance of the zero columns in the reduced partition. There are two ways to overcome this problem. The first one considered in Section 5.1 is quite straightforward and it just rewrites the system ${\bf s}={\bf D}\cdot{\bf x}$ transferring all C-columns to the l.h.s. so that the modified matrix can be reduced using the algorithms discussed in the preceding Sections.
The second approach is more elaborated as it is based on the extension of the original equations ${\bf s}={\bf D}\cdot{\bf x}$ to a larger but equivalent system $\tilde{\bf s}=\tilde{\bf D}\cdot\tilde{\bf x}$ where the matrix $\tilde{\bf D}$ is void of C- and NP-columns. The extended partition can be then reduced by the Cayley elimination procedure and we consider the details of this process in Section 5.2. Finally, in Section 5.3 we compare these two approaches.
5.1 Partition convolution superposition
Call an integer generator matrix competent if its columns (viewed as vectors) are noncollinear and incompetent otherwise. Assume that the vectors corresponding to the first $n$ columns of the incompetent generator matrix $\bf D$} are parallel to each other and rewrite the linear system ${\bf s}={\bf D}\cdot{\bf x}$ as follows